OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 12 cm and a length of 16 cm? A. P = 2 · 12 + 2 · 16 B. P = 2 · 12 + 16 C. P = 12 · 16 D. P = 12 + 16

OpenStudy (anonymous):

please help i will give a medal

OpenStudy (anonymous):

i was thinking b

OpenStudy (kaylaprincess):

p = l * w

OpenStudy (kaylaprincess):

what do you think it would be now?

OpenStudy (anonymous):

so c

OpenStudy (kaylaprincess):

correct :-)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its not c

OpenStudy (anonymous):

it's not

OpenStudy (kaylaprincess):

what do you say it is?

OpenStudy (anonymous):

its a

OpenStudy (kaylaprincess):

OH MY GOSH

OpenStudy (anonymous):

you little #####

OpenStudy (anonymous):

no wonder i failed my test

OpenStudy (anonymous):

I PROMISE IT IS A

OpenStudy (kaylaprincess):

my apologies ! I was not thinking for this question! thank you @polpolpol12

OpenStudy (kaylaprincess):

You failed your test because you didn't know the material. Why were you posting your tests questions? That is an integrity violation and you could be reported.

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

it a

OpenStudy (anonymous):

It is A because in there are two sides of length and height. So P=l+l+h+h or P=2*l+2*h l=length and h=height Hope this helps!

OpenStudy (anonymous):

first of all kaylaprincess i forgot it was really a assessment so really i should not be reported because it was not a test come to think about it it was only homework so...

OpenStudy (anonymous):

and also homework made to be somethings not even u know and everyone needs help at some point.Maybe not you because you think you know everything so you could shut your mouth and i will shut mine

OpenStudy (anonymous):

@cyborg88: If it's not an assessment, you've got nothing to worry about. In the future, for advice, I suggest you keep your anger contained and not to express it on OS. But I understand it is really annoying when somebody is telling you should be something or whatever. I feel you, so no hard feelings bro.